OpenStudy (anonymous):

I am terrible at math - please help! How do I add fractions with unlike denominators? such as 21/110 + 1/11 ???

OpenStudy (anonymous):

Find the LCM. (Least common multiple). That is 110. So you have to multiply the 11 in 1/11 by 10 in order to get 110. You have to multiply the top as well as the bottom. So it looks like this: 10/110 plus 21/110. 10 plus 21 is 31. Answer is 31/110. :)

OpenStudy (anonymous):

thank you

OpenStudy (thomas5267):

Use LaTeX! \[ \begin{align*} &\frac{21}{110}+\frac{1}{11} \\ =&\frac{21}{110}+\frac{1\times10}{11\times10} \\ =&\frac{21}{110}+\frac{10}{110} \\ =&\frac{31}{110} \end{align*} \]

OpenStudy (anonymous):

LaTex???

OpenStudy (anonymous):

No problem. :)

OpenStudy (thomas5267):

That thing that shows mathematical formula.

OpenStudy (anonymous):

OHHH! Thx